OpenStudy (anonymous):

when using distance formula to find the distance between two points, how would you write the answer if it isn't a perfect square?

OpenStudy (anonymous):

try and figure out if its a fraction if so use that if you cant find the fraction i would say round to nearest thousandth, unless specified that you need exact answer

OpenStudy (dumbcow):

Example: distance between (1,4) and (-2,1) \[d = \sqrt{(1-(-2))^{2} + (4-1)^{2}} = \sqrt{3^{2}+3^{2}} = \sqrt{2*3^{2}} = \sqrt{3^{2}}*\sqrt{2} = 3\sqrt{2}\]

OpenStudy (anonymous):

@dumbcow would you write your answer as 3 radical 2 0r as a decimal or fraction. Does it matter???

OpenStudy (dumbcow):

depends on the what they are asking for. Usually they will either ask for an exact answer or a decimal approximation exact -> 3sqrt(2) decimal -->4.24

OpenStudy (anonymous):

it just says to find the distance between the two points, so how would you write it?

OpenStudy (dumbcow):

i dunno either one works, what does your teacher usually want i would go with the exact answer

OpenStudy (anonymous):

i dunno my teacher yet. it's a summer packect but thanks :)